OpenStudy (king_of_kings):

Expand and evaluate the power. The third power of 2.5. A. 0.833 B. 5.5 C. 7.5 D. 15.625

OpenStudy (king_of_kings):

i need help

OpenStudy (phi):

are you allowed to use a calculator ?

OpenStudy (phi):

\[ 2.5^3 \] means 2.5*2.5*2.5 many calculators also have a "cube" button (raise to the 3rd power), so you could use that or you could write 2.5 as 5/2 and do \[ \frac{5}{2} \cdot \frac{5}{2} \cdot \frac{5}{2} = \frac{125}{8} \]

OpenStudy (phi):

based on the answer choices, you could also do this: 2 *2 *2 = 8 3*3*3= 27 and we know 2.5 is between 2 and 3 so 2.5*2.5 * 2.5 is between 8 and 27 and we see only 1 of the choices is between 8 and 27